OpenStudy (anonymous):

Which of the following is the correct mathematical expression for of entropy?

OpenStudy (anonymous):

A. ΔS = ΔV - TΔS B. ΔS = ΔQ - TΔS C. ΔS = ΔQ - TΔG D.

OpenStudy (sweetburger):

where is choice D?

OpenStudy (aaronq):

if you're trying to rearrange \(\Delta G =\Delta H -T\Delta S\), then it's \(\Delta S=\dfrac{\Delta H-\Delta G}{T}\)